(22-23) You borrow $100,000 Constant Amortization Mortgage (CAM) at 10% for 20 years, assuming annual payment, answer-

22. What is your total payment for year 11?

a. 5,000

b. 6,000

c. 8,000

d. 10,000

23. What is the total principle you paid in year 11?

a. 2,000

b. 3,000

c. 4,000

d. 5,000
Answer #1

In case of Constant Amortization Mortgage, the loan amount is divided in the years of loan period and each year's installment involves the part of loan divided in years plus the interest being charged on the remaining balance of loan.

Here, the loan amount is $100000 for 20 years

Each year installment will have 100000/20 = $ 5000 as the principle amount. Above this, interest for each year will be charged for the remaining portion of the principle amount.

Payment for year 11 :

Principle amount = $ 5000

We must have paid 10 installments, thus 10*5000 = 50000 of the principle amount has been paid

Remaining loan amount = 100000 - 50000 = $50000

Interest = 50000*10% = 5000

Thus total payment for year 11 = Principle + Interest = $5000 + $ 5000 = $10000

22. Correct answer is (d.) $10000

Now, the total principle paid in year 11 will be $ 5000 (as explained earlier)

Thus, correct answer for 23. is (d.) $ 5000